what is the sum of the measures of the vertex angles of a 12 sided polygon

Answer by RAY100(1637)

12 sides polygon
.
remember the sum of interior angles of a polygon are (n-2) * 180
{ triangle,,3 sided,,(3-2)*(180)=180 }
.
Individual interior angles are { (n-2) * 180} / n
{ triangle is 180 / 3 = 60}
.
For 12 sided polygon,,sum of interior angles is
.
{ (n-2) * (180) } = (12-2)(180) = 1800 degrees
.
individual interior angle is,,1800/12 = 150 degrees